# The 2-systole on compact Kähler surfaces with positive scalar curvature

### v1

@@ 4.3. Rational ruled surfaces
Setting $x:=a/t,\penalty 10000\ y:=b/t$, so that $a=xt$, $b=yt$, then the Kähler conditions $t>0$, at>0a-t>0, bt>0b-t>0, and $\omega]^{2}>0$ translate into
$x>1,\quad y>1,\quad\text{and }\penalty 10000\ \penalty 10000\ 2xy>1.$
Thus,
@@
defined on $\{(x,y)\mid x>1,\ y>1\}$. Observe that
$\sup_{x,y}\left(\min\{1,x-1,y-1\}\frac{2x+2y-1}{2xy-1}\right)=1$
on this domain, with equality at $x=y=2$. We then conclude
@@
A direct computation of the partial derivatives shows that $\phi_{e}$ has a unique critical point in $\mathcal{D}_{e}$ at $(x,z)=(2,1)$, so that this point realizes a global maximum. At $(x,z)=(2,1)$ one has
$\sup_{(x,z)\in\mathcal{D}_{e}}\phi_{e}(x,z)=\phi_{e}(2,1)=\frac{2e+5}{4e+3}.$
@@
Let $F_{1}=[\mathbb{P}^{1}\times\{\mathrm{pt}\}]$ and $F_{2}=[\{\mathrm{pt}\}\times\mathbb{P}^{1}]$ be the two rulings, and let $E_{1},\dots,E_{k}$ be the exceptional curves. Since the two rulings are algebraically equivalent and $\mathcal{J}_{X_{k}}$ is homogeneous of degree zero, we may assume that the coefficient of the “smaller” ruling is $1$. More precisely, any Kähler class can be written (after possibly interchanging $F_{1},F_{2}$ and rescaling) in the form $\omega]=F_{1}+bF_{2}-\sum t_{i}E_{i},$ for $b\geq 1,\ t_{i}>0$. Set $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. The effective curves $E_{i},\ F_{1}-E_{i},\ F_{2}-E_{i}$ on $X_{k}$ yield the Kähler condition
@@
Write any Kähler class on $X_{k}$ in the form $\omega]\;=\;C_{0}+bF-\sum t_{i}E_{i}$, for $b>e,\;t_{i}>0$. We use the same notations as before, so that
@@
with the inequality holds strictly if $e\geq 2$. In particular, $F_{e}(m,t)\geq 0$ for all admissible $m,t$, and hence $\phi_{e,k}(b,m,S)\leq 1$.
@@
Fix an integer $e\geq 1$ and let $X_{0}=\mathbb{F}_{e}$ be the $e$-th Hirzebruch surface with section $C_{0}$ and fibre class $F$. Let $X_{k}:=\text{Bl}_{k}(\mathbb{F}_{e})$ be the blow-up of $k$ points $p_{1},\dots,p_{k}$ with $k\leq e$ in very general position. In this range of parameters, the surface $X_{k}$ is anti-canonical and the Mori cone $\overline{\mathrm{NE}}(X_{k})$ is polyhedral. In particular, see for instance [HJNK25, Proposition 2.4 and Lemma 3.4], every extremal ray of $\overline{\mathrm{NE}}(X_{k})$ is generated by $C_{0},\penalty 10000\ E_{i},\penalty 10000\ F_{i}-E_{i}$ for $1\leq i\leq k$. We now fix a Kähler class on $X_{k}$ and normalize it as in Lemma 4.8:
@@
$m:=\operatorname{sys}_{2}(\omega)=\min_{1\leq i\leq k}\bigl\{t_{i},1-t_{i},b-e\bigr\}.$
From $t_{i}\geq m$ and $1-t_{i}\geq m$ we immediately deduce $0<m\leq\frac{1}{2}$. Therefore the functional $\mathcal{J}_{X_{k}}([\omega])$ takes the explicit form
@@
and the variables satisfy the Kähler inequalities
@@
In this regime, we can specifically calculate the maximum of the smooth function
$(b,t_{1},\dots,t_{k})\longmapsto m\cdot\frac{2b+2-e-S}{2b-e-Q}$
under the above constraints. In this case, the maximum is achieved at (e+1/2,1/2,,1/2)(e+1/2,1/2,...,1/2), so that
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=4\pi\penalty 10000\ \frac{2e+6-k}{4e+4-k}\leq 4\pi$
for all such Kähler classes.
@@
Moreover, equality holds if and only if $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ (equivalently, the minimal model is $\mathbb{F}_{0}$), endowed with the product Fubini-Study metric.
@@ 5. Level set method on non-rational PSC Kähler surfaces
In [Ste22], Stern introduced the following inequality for a non-constant $S^{1}$-valued harmonic map $u$ on a $3$-manifold $(M,g)$ through the level set method:
@@ 5.1. A Stern-type scalar curvature inequality
Since $D=D_{z}$ is a fibre of the holomorphic fibration $f\colon X\to B$, its normal bundle $\mathcal{N}_{D}\cong\mathcal{O}(D)|_{D}$ is holomorphically trivial. Taking the first Chern class in the adjunction formula, we get
$c_{1}(D)=c_{1}(X)|_{D},$
and hence, for the Ricci forms with respect to $\omega$,
$\operatorname{Ric}_{D}(\omega)=\operatorname{Ric}_{X}(\omega)|_{D}.$
Let $\nu$ be a local unit normal vector field of type $(1,0)$ along $D$ with respect to $\omega$. Tracing the Gauss equation yields
$S_{D}(\omega)=\operatorname{tr}_{\omega}\operatorname{Ric}_{D}(\omega)=\operatorname{tr}_{\omega}\operatorname{Ric}_{X}(\omega)|_{D}=S_{X}(\omega)-\operatorname{Ric}_{X}(\omega)(\nu,\bar{\nu}),$
where $S_{X}(\omega)$ and $S_{D}(\omega)$ denote the scalar curvatures of $(X,\omega)$ and $(D,\omega|_{D})$, respectively. In particular, since $\nu=\nabla^{1,0}f/|\nabla^{1,0}f|$, we obtain
$\operatorname{Ric}_{X}(\omega)\bigl(\nabla^{1,0}f,\nabla^{0,1}f\bigr)=\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr).$ (5.2)
@@
Let $f\colon X\to B$ be a non-constant holomorphic map, and let $g\in C^{\infty}(X)$. Then for every regular value $z\in B$ of $f$ we have
@@
$\Delta|\partial f|^{2}\;\geq\;|\nabla\partial f|^{2}+|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr).$ (5.3)
@@
Let $(X^{n},\omega)$ be a compact Kähler manifold and let $(B,\omega_{0})$ be a compact Riemann surface of genus $g(B)\geq 1$, endowed with a constant curvature metric $\omega_{0}$. Suppose that $f\colon X\rightarrow B$ is a non-constant holomorphic map, and let $D=f^{-1}(z)$ denote a regular fibre. Then for any $\phi\in C^{\infty}(X)$ we have
$\int_{B}\left[\int_{D}\phi^{2}\left(\frac{\left|\nabla\partial f\right|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0}\;\leq\;-\,n\int_{X}\sqrt{-1}\,\partial(\phi^{2})\wedge\bar{\partial}|\partial f|^{2}\wedge\omega^{n-1}.$ (5.4)
@@
Write $B=A\cup B_{0}$, where $A$ is an open neighbourhood of the (finite) set of critical values of $f$ and $B_{0}$ consists only of regular values. Multiplying (5.3) by $\phi^{2}$ and integrating over $f^{-1}(B_{0})$, we obtain
$\int_{f^{-1}(B_{0})}\phi^{2}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr)\Bigr)\omega^{n}\;\leq\;\int_{f^{-1}(B_{0})}\phi^{2}\bigl(\Delta|\partial f|^{2}\bigr)\omega^{n}.$ (5.5)
For the right-hand side of (5.5), integration by parts gives
$\int_{f^{-1}(B_{0})}\phi^{2}\bigl(\Delta|\partial f|^{2}\bigr)\,\omega^{n}=n\int_{f^{-1}(B_{0})}\phi^{2}\sqrt{-1}\,\partial\bar{\partial}|\partial f|^{2}\wedge\omega^{n-1}.$
On the other hand, applying the co-area formula to the left-hand side of (5.5) yields
$\int_{f^{-1}(B_{0})}$ $\phi^{2}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}-S_{D}\bigr)\Bigr)\omega^{n}$
$=n\int_{B_{0}}\left[\int_{D}\phi^{2}\left(\frac{\left|\nabla\partial f\right|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0}.$
By Sard’s theorem, the set of critical values has measure zero, and by choosing $A$ with arbitrarily small measure we can pass to the limit and replace $f^{-1}(B_{0})$ and $B_{0}$ by $X$ and $B$, respectively. Finally, another integration by parts shows that
$n\int_{X}\phi^{2}\sqrt{-1}\,\partial\bar{\partial}|\partial f|^{2}\wedge\omega^{n-1}=-\,n\int_{X}\sqrt{-1}\,\partial(\phi^{2})\wedge\bar{\partial}|\partial f|^{2}\wedge\omega^{n-1},$
and combining the above identities with (5.5) gives (5.4). The equality statement follows from the discussion before (5.3): equality in the Bochner formula holds if and only if $\operatorname{Rm}(\omega_{0})\equiv 0$, i.e. $g(B)=1$ and $(B,\omega_{0})$ is a flat elliptic curve. ∎

### v2

@@ 4.3. Rational ruled surfaces
Setting $x:=a/t,\penalty 10000\ y:=b/t$, so that $a=xt$, $b=yt$, then the Kähler conditions $t>0$, at>0a-t>0, bt>0b-t>0, and $\omega]^{2}>0$ translate into $x>1,\quad y>1,$ and $2xy>1$. Thus,
@@
defined on $\{(x,y)\mid x>1,\ y>1\}$. A straightforward piecewise analysis of the two-variable function
$(x,y)\longmapsto\min\{1,x-1,y-1\}\frac{2x+2y-1}{2xy-1},\qquad x>1,\ y>1,$
shows that its supremum is $1$, attained uniquely at $x=y=2$. We then conclude
@@
The function $\phi_{e}(x,z)$ is piecewise rational, according to the regions where $\min\{1,x-1,z\}$ is realized by one of its three entries. A straightforward inspection on these regions shows that $\phi_{e}$ attains its global maximum at $(x,z)=(2,1)$, with value
$\phi_{e}(2,1)=\frac{2e+5}{4e+3}.$
@@
Let $F_{1}=[\mathbb{P}^{1}\times\{\mathrm{pt}\}]$ and $F_{2}=[\{\mathrm{pt}\}\times\mathbb{P}^{1}]$ be the two rulings, and let $E_{1},\dots,E_{k}$ be the exceptional curves. By the symmetry between the two rulings and the scale invariance of $\mathcal{J}_{X_{k}}$, we may, after possibly interchanging $F_{1}$ and $F_{2}$ and rescaling, assume that the coefficient of the smaller ruling is $1$. More precisely, any Kähler class can be written (after possibly interchanging $F_{1},F_{2}$ and rescaling) in the form $\omega]=F_{1}+bF_{2}-\sum t_{i}E_{i},$ for $b\geq 1,\ t_{i}>0$. Set $m=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$. The effective curves $E_{i},\ F_{1}-E_{i},\ F_{2}-E_{i}$ give candidates for $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$, hence
@@
After rescaling, write any Kähler class on $X_{k}$ in the form $\omega]\;=\;C_{0}+bF-\sum t_{i}E_{i}$, for $b>e,\;t_{i}>0$. We use the same notations as before, so that
@@
with strict inequality if $e\geq 2$. In particular, $F_{e}(m,z)\geq 0$ for all admissible $m,z$, and hence $\phi_{e,k}(b,m,S)\leq 1$.
@@
Fix an integer $e\geq 1$ and let $X_{0}=\mathbb{F}_{e}$ be the $e$-th Hirzebruch surface with section $C_{0}$ and fibre class $F$. Let $X_{k}:=\text{Bl}_{k}(\mathbb{F}_{e})$ be the blow-up of $k$ points $p_{1},\dots,p_{k}$ with $k\leq e$ in very general position. In this setting, the Mori cone $\overline{\mathrm{NE}}(X_{k})$ is polyhedral. Moreover, see for instance [13, Proposition 2.4 and Lemma 3.4], every extremal ray of $\overline{\mathrm{NE}}(X_{k})$ is generated by $C_{0},\penalty 10000\ E_{i},\penalty 10000\ F-E_{i}$ for $1\leq i\leq k$. We now fix a Kähler class on $X_{k}$ and normalize it as in Lemma 4.8:
@@
$m:=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])=\min\bigl\{t_{i},\ 1-t_{i},\ b-e:\ 1\leq i\leq k\bigr\}.$
From $t_{i}\geq m$ and $1-t_{i}\geq m$ we immediately deduce $0<m\leq\frac{1}{2}$. The functional $\mathcal{J}_{X_{k}}([\omega])$ takes the explicit form
@@
and the parameters satisfy
@@
Assume in addition that $k\leq 8$. Then a direct optimization of the above function under these constraints shows that the maximum is achieved at (e+1/2,1/2,,1/2)(e+1/2,1/2,...,1/2). Consequently,
$\sup_{\mathcal{K}^{+}(X_{k})}\mathcal{J}_{X_{k}}([\omega])=4\pi\,\frac{2e+6-k}{4e+4-k}\leq 4\pi,\qquad k\leq\min\{e,8\}.$
Remark 4.10.
The threshold $k=8$ in Example 4.9 has a genuine structural meaning. From the algebro-geometric point of view, since $K_{\mathbb{F}_{e}}^{2}=8$ and each blow-up decreases $K^{2}$ by $1$, one has $K_{X_{k}}^{2}=8-k$. Thus $k=8$ is precisely the point where the sign of $K_{X_{k}}^{2}$ changes. This is the analogue, for blow-ups of Hirzebruch surfaces, of the classical transition at $k=9$ for blow-ups of $\mathbb{P}^{2}$.
The same threshold also appears directly in the optimization of $\mathcal{J}_{X_{k}}$. Indeed, writing
$\Phi(b,t):=m\,\frac{2b+2-e-S}{2b-e-Q},\qquad m=\min\{t_{i},1-t_{i},b-e\},$
one computes
$\partial_{b}\Phi=\frac{2m(S-Q-2)}{(2b-e-Q)^{2}},\qquad S-Q=\sum_{i}t_{i}(1-t_{i})\leq\frac{k}{4}.$
[…]
@@
Moreover, equality holds if and only if $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$, endowed with the product Fubini-Study metric.
Proof.
Let $X\to\mathbb{P}^{1}$ be a rational ruled surface endowed with a PSC Kähler metric $\omega$. Since every rational ruled surface is obtained from a Hirzebruch surface $\mathbb{F}_{e}$ by blowing up finitely many points, we distinguish cases according to the number of blow-ups.
If $X$ is minimal, then $X\cong\mathbb{F}_{e}$ for some $e\geq 0$. By Proposition 4.2,
$\sup_{\mathcal{K}(\mathbb{F}_{e})}\mathcal{J}_{\mathbb{F}_{e}}([\omega])=4\pi\,\frac{e+4}{e+2}\leq 8\pi,$
with equality if and only if $e=0$, i.e. $X\cong\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$.
If $X$ is the blow-up of $\mathbb{F}_{e}$ at one point, then Proposition 4.6 gives
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])=\begin{cases}4\pi,&e=0,\\[4.0pt] 4\pi\,\dfrac{2e+5}{4e+3},&e\geq 1,\end{cases}$
and in either case this is strictly smaller than $8\pi$.
If $X$ is the blow-up of $\mathbb{F}_{e}$ at $k\geq 2$ points, then Lemma 4.7 and 4.8 gives
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])\leq 4\pi.$
Combining all cases, we obtain
$\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}([\omega])\leq 8\pi,$
and equality can occur only when $X\cong\mathbb{F}_{0}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. Now apply Theorem 2.2 (1). Since $\omega$ is PSC, we have $\omega]\in\mathcal{K}^{+}(X)$, and therefore
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)\leq\mathcal{J}_{X}([\omega])\leq 8\pi.$
This proves the inequality.
It remains to characterize the equality case. Suppose $\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)=8\pi$. Then necessarily $\mathcal{J}_{X}([\omega])=8\pi$, so by the discussion above we must have $X\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$. Moreover, equality in Theorem 2.2 (1) implies that $\omega$ is cscK and that $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic curve. On $\mathbb{P}^{1}\times\mathbb{P}^{1}$, write $\omega]=aF_{1}+bF_{2}$, for $a,b>0$ where $F_{1},F_{2}$ are the two rulings. Assuming without loss of generality that $a\leq b$, then
$\mathcal{J}_{\mathbb{P}^{1}\times\mathbb{P}^{1}}([\omega])=4\pi\,a\,\frac{a+b}{ab}=4\pi\left(1+\frac{a}{b}\right)\leq 8\pi,$
with equality if and only if $a=b$. Thus $\omega]$ is proportional to $F_{1}+F_{2}$. Since $\omega$ is cscK, it follows that, up to overall scaling, $\omega$ is the product Fubini–Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
Conversely, the product Fubini-Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ has constant scalar curvature, and its $2$-systole is realized by either ruling $\mathbb{P}^{1}$. A direct computation shows that equality holds:
$\operatorname{sys}_{2}(\omega)\cdot\min_{X}S(\omega)=8\pi.$
This completes the proof. ∎
@@ 5. Level set method on non-rational PSC Kähler surfaces
In [22], Stern introduced the following inequality for a non-constant $S^{1}$-valued harmonic map $u$ on a $3$-manifold $(M,g)$ through the level set method:
@@ 5.1. A Stern-type scalar curvature inequality
Since $D=D_{z}$ is a fibre of the holomorphic fibration $f\colon X\to B$, its normal bundle $\mathcal{N}_{D}\cong\mathcal{O}(D)|_{D}$ is holomorphically trivial. By the adjunction formula,
$K_{D}=\bigl(K_{X}\otimes\mathcal{O}(D)\bigr)|_{D}.$
Taking first Chern classes, we obtain
$c_{1}(D)=c_{1}(X)|_{D}.$
Therefore, for the Ricci forms induced by $\omega$, the difference
$\operatorname{Ric}_{D}(\omega)-\operatorname{Ric}_{X}(\omega)|_{D}$
is an exact real $(1,1)$-form on the compact Riemann surface $D$. By the $\partial\bar{\partial}$-lemma, there exists a smooth real-valued function $\phi\in C^{\infty}(D)$ such that
$\operatorname{Ric}_{D}(\omega)=\operatorname{Ric}_{X}(\omega)|_{D}+\sqrt{-1}\,\partial_{D}\bar{\partial}_{D}\phi.$
Tracing with respect to the induced metric $\omega|_{D}$, we get
$S_{D}(\omega)=\operatorname{tr}_{\omega|_{D}}\bigl(\operatorname{Ric}_{X}(\omega)|_{D}\bigr)+\Delta_{D}\phi.$
Let $\nu$ be a local unit normal vector field of type $(1,0)$ along $D$ with respect to $\omega$. Since $D\subset X$ is a complex hypersurface, we have
$\operatorname{tr}_{\omega|_{D}}\bigl(\operatorname{Ric}_{X}(\omega)|_{D}\bigr)=S_{X}(\omega)-\operatorname{Ric}_{X}(\omega)(\nu,\bar{\nu}).$
Hence
$S_{D}(\omega)=S_{X}(\omega)-\operatorname{Ric}_{X}(\omega)(\nu,\bar{\nu})+\Delta_{D}\phi.$
In particular, since
$\nu=\frac{\nabla^{1,0}f}{|\nabla^{1,0}f|},$
we obtain
$\operatorname{Ric}_{X}(\omega)\bigl(\nabla^{1,0}f,\nabla^{0,1}f\bigr)=|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr).$ (5.2)
@@
Let $f\colon X\to B$ be a non-constant holomorphic map, and let $g\in C^{\infty}(X)$. Then
@@
$\Delta|\partial f|^{2}\;\geq\;|\nabla\partial f|^{2}+|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr).$ (5.3)
@@
Let $(X^{n},\omega)$ be a compact Kähler manifold and let $(B,\omega_{0})$ be a compact Riemann surface of genus $g(B)\geq 1$, endowed with a constant curvature metric $\omega_{0}$. Suppose that $f\colon X\rightarrow B$ is a non-constant holomorphic map, and let $D=f^{-1}(z)$ denote a regular fibre. Then
$\int_{B}\left[\int_{D}\left(\frac{\left|\nabla\partial f\right|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0}\;\leq\;0.$ (5.4)
@@
Write $B=A\cup B_{0}$, where $A$ is an open neighbourhood of the finite set of critical values of $f$, and $B_{0}$ consists only of regular values. Integrating (5.3) over $f^{-1}(B_{0})$, we obtain
$\int_{f^{-1}(B_{0})}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr)\Bigr)\omega^{n}\;\leq\;\int_{f^{-1}(B_{0})}\bigl(\Delta|\partial f|^{2}\bigr)\omega^{n}.$ (5.5)
Since $X$ is compact and has no boundary,
$\int_{X}\bigl(\Delta|\partial f|^{2}\bigr)\omega^{n}=0.$
Hence, by choosing $A$ with arbitrarily small measure and using Sard’s theorem, we may pass to the limit in (5.5) and obtain
$\int_{X}\Bigl(\bigl|\nabla\partial f\bigr|^{2}+\bigl|\nabla^{1,0}f\bigr|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)+\Delta_{D}\phi\bigr)\Bigr)\omega^{n}\;\leq\;0.$
Now apply the co-area formula term by term. Since $|\nabla^{1,0}f|^{2}=|\partial f|^{2}$, we get
$\int_{X}|\nabla\partial f|^{2}\,\omega^{n}$ $=n\int_{B}\left(\int_{D}\frac{|\nabla\partial f|^{2}}{|\partial f|^{2}}\,\omega^{n-1}\right)\omega_{0},$
$\int_{X}|\nabla^{1,0}f|^{2}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr)\,\omega^{n}$ $=n\int_{B}\left(\int_{D}\bigl(S_{X}(\omega)-S_{D}(\omega)\bigr)\,\omega^{n-1}\right)\omega_{0},$
and similarly
$\int_{X}|\nabla^{1,0}f|^{2}\,\Delta_{D}\phi\,\omega^{n}=n\int_{B}\left(\int_{D}\Delta_{D}\phi\,\omega^{n-1}\right)\omega_{0}.$
Since each regular fibre $D$ is compact without boundary,
$\int_{D}\Delta_{D}\phi\,\omega^{n-1}=0.$
Therefore
$0$ $\geq n\int_{B}\left[\int_{D}\left(\frac{|\nabla\partial f|^{2}}{|\partial f|^{2}}+S_{X}(\omega)-S_{D}(\omega)\right)\omega^{n-1}\right]\omega_{0},$
which gives (5.4). The equality statement follows from the discussion before (5.3): equality in the Bochner inequality holds if and only if $\operatorname{Rm}(\omega_{0})\equiv 0$, namely if and only if $g(B)=1$ and $(B,\omega_{0})$ is flat. ∎